OpenStudy (anonymous):

What number comes next in this pattern? 22, 26, 30, 34, 38, 42, ___

OpenStudy (anonymous):

@CausticSyndicalist

OpenStudy (anonymous):

@acxbox22

OpenStudy (haleyelizabeth2017):

What is 14+52?

OpenStudy (anonymous):

66

OpenStudy (anonymous):

*sigh*

OpenStudy (haleyelizabeth2017):

There ya go! That's your answer :P

OpenStudy (acxbox22):

is there supposed to be an x in front of -52 if it is like x-52=14 if so then add 52 on both sides which leaves x=66

OpenStudy (anonymous):

What number could replace the box to make this number sentence true? 35 – 17 = 11 + A. 22 B. 18 C. 8 D. 7

OpenStudy (haleyelizabeth2017):

so first, you need to subtract the first two terms. What do you get?

OpenStudy (haleyelizabeth2017):

So here are the steps. (I put x where the box is :P) 35-17=18 18=11+x 18-11=x x=7

OpenStudy (acxbox22):

please post each question separately so you can receive maximum help :) thank you

OpenStudy (anonymous):

@sammixboo